OpenStudy (anonymous):

In triangle ABC angle a measures twice angle b. If an exterior angle at C measures 126, find angle a

jhonyy9 (jhonyy9):

than C =180 - 126 =54 degree so than 180=a+b+c 180=2b+b+54 126=3b b=126/3 =42 b=42 a=2b=2*42=84 a=84 b=42 c=54